|x-1|
-----=|1-x|+5
-3

Answer by stanbon(75887)

|x-1|
-----=|1-x|+5
-3
====================
|x-1| = -3|1-x| -15
---
But |1-x| = |x-1|
----
So, 4|x-1| = -15
---
No solution since an absolute value cannot be negative.
-------------
Cheers,
Stan H.